# Accretion of Generalized Chaplygin Gas onto Cosmologically Coupled Black Holes

### v1

@@ IV.1 Dark energy accretion
A similar equation appears in Ref. [babichev2012backreaction], but in the case where $H_{0}=0$. In Ref. [babichev2012backreaction] treatment, Einstein’s equations are evaluated in a vacuum background, that is, $G_{\mu\gamma}\left[g_{\mu\gamma}^{(0)}\right]=0$. In contrast, our zeroth-order solution to Einstein’s equations is given by the McVittie equations.
IV.2 Generalized Chaplygin Gas
@@
Furthermore, there exists an effective equation of state in the intermediate regime between the matter-dominated and de Sitter phases, which can be obtained by expanding Eq. (7) as in Ref. [bento2002generalized]:
$\rho_{GCG}$ $\approx$ $A^{1/\left(1+\alpha\right)}+\left(\frac{1}{1+\alpha}\right)\frac{B}{A^{\alpha/\left(1+\alpha\right)}}a^{-3\left(1+\alpha\right)}+$...,
$p_{GCG}$ $\approx$ $-A^{1/\left(1+\alpha\right)}+\left(\frac{\alpha}{1+\alpha}\right)\frac{B}{A^{\alpha/\left(1+\alpha\right)}}a^{-3\left(1+\alpha\right)}+$...
Eqs. (47) are essential, since within the perturbative framework we are using, the contribution of the accreting fluid must be much smaller than those of the background fluid.
However, since we are assuming that the GCG drives the cosmic expansion, thus determining the background metric, while we are also interested in analyzing its accretion, we shall adopt the following strategy: we associate the dominant term in Eq. (47) with the background fluid characterized by $H_{0}$, and the remaining term with the perturbation.
@@ IV.2.1 Matter dominated era
According to our perturbative construction, we shall assume that the accreting fluid obeys the subdominant term of the equation of state given in Eq. (47). Thus, its equation of state reads
@@
$F\left(t\right)$ $=$ $0,$ (52a)
$H_{1}^{2}\left(t\right)$ $=$ $\frac{4}{9t^{2}}+\frac{8\pi A^{1/\left(1+\alpha\right)}}{3},$ (52b)
$M_{1}\left(t\right)$ $=$ $M_{0}+M_{ac}+\left(H_{0}^{2}\left(t_{0ac}\right)-H_{1}^{2}\left(t\right)\right)\frac{r_{0ac}^{3}}{2},$
@@
We now study how accretion affects the evolution of the cosmological and black hole apparent horizons in comparison with the non-accreting case. To this end, we compare Eqs. (12a) and (12) with Eqs. (39).
From Eqs. (12a) and (12), as discussed in section II, when the equation
@@
Analogously, we can identify the instant $t_{0ac}$ corresponding to the formation of the horizons due to accretion. Similarly to the non-accreting case, from Eqs. (39) it follows that the horizons are formed at instant $t_{0ac}$ satisfying
@@
Using Eqs. (52b) and (LABEL:M1_matter), Eq. (57) admits, as its positive solution,
@@
Figure 1: Apparent black hole and cosmological horizon formation time as a function of AA (solid lines), as given by (58). The dashed line represent the instant of formation of the apparent horizons without available matter for accretion, $t_{0}$, given by 56. The mass parameter $M_{0}$ is set to 11. We also consider $\alpha=0.1$.
@@
The horizons obtained from Eqs. (39), together with Eqs. (52b) and (LABEL:M1_matter), are shown in Fig. 2 for different values of $A$. As illustrated, larger values of $A$, and therefore higher energy densities of matter available to the accretion process, cause the black hole horizon to expand, starting from an initial radius of $3M_{0}$ and settling at constant values greater than the Schwarzschild radius $2M_{0}$ after a sufficiently long time. In contrast, the cosmological horizon decreases with increasing $A$, eventually reaching a steady monotonic regime.
@@
Let us now examine the horizons in their asymptotic stage, beginning with the case without accretion. After a sufficiently long time, such that $t>>t_{0ac}$, the horizons given by Eqs. (12a) and (12) behave as
$r_{b}^{\left(0\right)}\left(t\right)\approx 2M_{0}\text{, \ }r_{c}^{\left(0\right)}\left(t\right)\approx\frac{1}{H_{0}}-M_{0}\approx\frac{3t}{2}-M_{0}.$ (59)
Similarly, for Eqs. (39), we find that for time scales such that $t>>t_{0ac}$, the behavior of the horizons in the presence of accretion is given by
$r_{b}^{\left(1\right)}\left(t\right)$ $\approx$ $2M_{1}\left(t\right)\approx 3M_{0}-36\pi A^{1/(1+\alpha)}M_{0}^{3},$ (60a)
$r_{c}^{\left(1\right)}\left(t\right)$ $\approx$ $\frac{1}{H_{1}\left(t\right)}-M_{1}\left(t\right)\approx\left(\frac{4}{9t^{2}}+\frac{4\pi A^{1/(1+\alpha)}}{3}\right)^{-1/2}$ (60b)
$-\frac{3M_{0}}{2}+18\pi A^{1/(1+\alpha)}M_{0}^{3},$
@@
For the black hole horizon with accretion to be larger than in the case without accretion, the condition $A<\left(12\pi M_{0}^{2}\right)^{-\left(1+\alpha\right)}$ must be satisfied, according to Eq. (60a). However, due to the constraint previously imposed by Eq. (58), this requirement is automatically fulfilled.
@@
Furthermore, it is worth noting that the values of the Chaplygin parameter $A$ explored here are chosen for illustrative clarity and exceed those expected from cosmological considerations. Observationally consistent models typically give $10^{-60}{$\mathrm{m}$}^{-2(1+\alpha)}<A<10^{-54}{$\mathrm{m}$}^{-2(1+\alpha)}$ for $\alpha\leq 0.1$, as in Refs. [aurich2018concordance, barreiro2008wmap]. Using enhanced values in our simulations makes the accretion effects numerically visible while preserving the qualitative behavior of the horizon evolution. The physically motivated range of $A$ would simply place these effects at correspondingly smaller amplitudes, without modifying the trends highlighted by our results.
@@
For a cosmologically coupled black hole like the McVittie solution, the backreaction of accretion on the metric alters this behavior. The inflow of accreting matter locally increases the energy density and consequently reduces the rate at which cosmic expansion decelerates, as shown in Eqs. (60a) and (52b). The resulting higher effective expansion rate counteracts the growth of the local gravitational term, namely, the apparent black hole horizon, as expressed in Eqs. (60a) and (LABEL:M1_matter), thereby delaying the instant at which the apparent horizon condition is satisfied.
IV.2.2 de Sitter era
@@
According to our perturbative approach, since the background is determined by an equation of state of the form $\rho^{(dS)}_{GCG}=-p^{(dS)}_{GCG}=A^{1/\left(1+\alpha\right)}$, we assume that the accreting fluid corresponds to the subdominant term in Eq. (47), namely
@@
As discussed previously, Refs. [aurich2018concordance, barreiro2008wmap] typically give $10^{-60}{$\mathrm{m}$}^{-2(1+\alpha)}<A<10^{-54}{$\mathrm{m}$}^{-2(1+\alpha)}$ for $\alpha\leq 0.1$. If we conservatively assume an initial accretion velocity $v_{0}=0.01c_{s}$, where $c_{s}=\sqrt{{\partial p}/{\partial\rho}}$ is the sound speed derived from the equation of state (64), we obtain $v_{0}=0.01\alpha^{1/2}$.
@@
$F\left(t\right)$ $=$ $64\pi H_{A}^{2}M_{0}^{4}(1+\alpha)\int_{t_{0ac}}^{t}\rho_{ac}d\bar{t}=$ (66)
$=$ $\frac{64\pi H_{A}M_{0}^{4}}{3}\left(\rho_{ac0}-\rho_{ac}\left(t\right)\right)$
@@
$M_{1}\left(t\right)$ $=$ $M_{0}+\frac{32\pi M^{3}_{0}}{3}\rho_{ac0}+18\pi\alpha M_{0}^{3}\rho_{ac}$ (71)
$+\frac{64H_{A}M_{0}^{4}\pi}{3}\left(\rho_{ac0}-\rho_{ac}\left(t\right)\right),$
@@ V Conclusions
Acknowledgements
The work of O.B. is partially supported by FCT (Fundação para a Ciência e Tecnologia, Portugal) through the project 2024.00252.CERN. The work of L.F.R. is supported by FCT through the grant 2024.06413.BD. Both authors are partially supported by UID/04650 - Centro de Física das Universidades do Minho e do Porto. M. C. B. thanks the University of Porto for hosting part of this study.

### v2

@@ 4.1 Dark energy accretion
A similar equation appears in Ref. [42], but in the case where $H_{0}=0$. In Ref. [42] treatment, Einstein’s equations are evaluated in a vacuum background, that is, $G_{\mu\gamma}\left[g_{\mu\gamma}^{(0)}\right]=0$. In contrast, our zeroth-order solution to Einstein’s equations is given by the McVittie equations.
4.2 Generalized Chaplygin Gas
@@
Furthermore, there exists an effective equation of state in the intermediate regime between the matter-dominated and de Sitter phases, which can be obtained by expanding Eq. (8) as in Ref. [10]:
$\rho_{GCG}$ $\approx A^{1/\left(1+\alpha\right)}+\left(\frac{1}{1+\alpha}\right)\frac{B}{A^{\alpha/\left(1+\alpha\right)}}a^{-3\left(1+\alpha\right)}+$..., (56)
$p_{GCG}$ $\approx-A^{1/\left(1+\alpha\right)}+\left(\frac{\alpha}{1+\alpha}\right)\frac{B}{A^{\alpha/\left(1+\alpha\right)}}a^{-3\left(1+\alpha\right)}+$... (57)
Eqs. (56)-(57) are essential, since within the perturbative framework we are using, the contribution of the accreting fluid must be much smaller than those of the background fluid.
However, since we are assuming that the GCG drives the cosmic expansion, thus determining the background metric, while we are also interested in analyzing its accretion, we shall adopt the following strategy: we associate the dominant term in Eq. (56)-(57) with the background fluid characterized by $H_{0}$, and the remaining term with the perturbation.
@@ 4.2.1 Matter-dominated era
According to our perturbative construction, we shall assume that the accreting fluid obeys the subdominant term of the equation of state given in Eq. (56)-(57). Thus, its equation of state reads
@@
$F\left(t\right)$ $=0,$ (62)
$H_{1}^{2}\left(t\right)$ $=\frac{4}{9t^{2}}+\frac{8\pi A^{1/\left(1+\alpha\right)}}{3},$ (63)
$M_{1}\left(t\right)$ $=M_{0}+M_{ac}+\left(H_{0}^{2}\left(t_{0ac}\right)-H_{1}^{2}\left(t\right)\right)\frac{r_{0ac}^{3}}{2},$ (64)
@@
We now study how accretion affects the evolution of the cosmological and black hole apparent horizons in comparison with the non-accreting case. To this end, we compare Eqs. (14) and (2) with Eqs. (LABEL:r1b)-(48).
From Eqs. (14) and (2), as discussed in section 2, when the equation
@@
Analogously, we can identify the instant $t_{0ac}$ corresponding to the formation of the horizons due to accretion. Similarly to the non-accreting case, from Eqs. (LABEL:r1b)-(48) it follows that the horizons are formed at instant $t_{0ac}$ satisfying
@@
Using Eqs. (63) and (64), Eq. (69) admits, as its positive solution,
@@
Figure 1: Apparent black hole and cosmological horizon formation time as a function of AA (solid lines), as given by (70). The dashed line represent the instant of formation of the apparent horizons without available matter for accretion, $t_{0}$, given by 68. The mass parameter $M_{0}$ is set to 11. We also consider $\alpha=0.1$.
@@
The horizons obtained from Eqs. (LABEL:r1b)-(48), together with Eqs. (63) and (64), are shown in Fig. 2 for different values of $A$. As illustrated, larger values of $A$, and therefore higher energy densities of matter available to the accretion process, cause the black hole horizon to expand, starting from an initial radius of $3M_{0}$ and settling at constant values greater than the Schwarzschild radius $2M_{0}$ after a sufficiently long time. In contrast, the cosmological horizon decreases with increasing $A$, eventually reaching a steady monotonic regime.
@@
Let us now examine the horizons in their asymptotic stage, beginning with the case without accretion. After a sufficiently long time, such that $t>>t_{0ac}$, the horizons given by Eqs. (14) and (2) behave as
$r_{b}^{\left(0\right)}\left(t\right)$ $\approx 2M_{0}$ (71)
$r_{c}^{\left(0\right)}\left(t\right)$ $\approx\frac{1}{H_{0}}-M_{0}\approx\frac{3t}{2}-M_{0}.$ (72)
Similarly, for Eqs. (LABEL:r1b)-(48), we find that for time scales such that $t>>t_{0ac}$, the behavior of the horizons in the presence of accretion is given by
$r_{b}^{\left(1\right)}\left(t\right)$ $\approx 2M_{1}\left(t\right)\approx 3M_{0}-36\pi A^{1/(1+\alpha)}M_{0}^{3},$ (73)
$r_{c}^{\left(1\right)}\left(t\right)$ $\approx\frac{1}{H_{1}\left(t\right)}-M_{1}\left(t\right)\approx\left(\frac{4}{9t^{2}}+\frac{4\pi A^{1/(1+\alpha)}}{3}\right)^{-1/2}-\frac{3M_{0}}{2}$ (74)
$+18\pi A^{1/(1+\alpha)}M_{0}^{3},$
@@
For the black hole horizon with accretion to be larger than in the case without accretion, the condition $A<\left(12\pi M_{0}^{2}\right)^{-\left(1+\alpha\right)}$ must be satisfied, according to Eq. (73). However, due to the constraint previously imposed by Eq. (70), this requirement is automatically fulfilled.
@@
Furthermore, it is worth noting that the values of the Chaplygin parameter $A$ explored here are chosen for illustrative clarity and exceed those expected from cosmological considerations. Observationally consistent models typically give $10^{-60}{$\mathrm{m}$}^{-2(1+\alpha)}<A<10^{-54}{$\mathrm{m}$}^{-2(1+\alpha)}$ for $\alpha\leq 0.1$, as in Refs. [12, 13]. Using enhanced values in our simulations makes the accretion effects numerically visible while preserving the qualitative behavior of the horizon evolution. The physically motivated range of $A$ would simply place these effects at correspondingly smaller amplitudes, without modifying the trends highlighted by our results.
@@
For a cosmologically coupled black hole like the McVittie solution, the backreaction of accretion on the metric alters this behavior. The inflow of accreting matter locally increases the energy density and consequently reduces the rate at which cosmic expansion decelerates, as shown in Eqs. (73) and (63). The resulting higher effective expansion rate counteracts the growth of the local gravitational term, namely, the apparent black hole horizon, as expressed in Eqs. (73) and (64), thereby delaying the instant at which the apparent horizon condition is satisfied.
4.2.2 de Sitter era
@@
According to our perturbative approach, since the background is determined by an equation of state of the form $\rho^{(dS)}_{GCG}=-p^{(dS)}_{GCG}=A^{1/\left(1+\alpha\right)}$, we assume that the accreting fluid corresponds to the subdominant term in Eq. (56)-(57), namely
@@
As discussed previously, Refs. [12, 13] typically give $10^{-60}{$\mathrm{m}$}^{-2(1+\alpha)}<A<10^{-54}{$\mathrm{m}$}^{-2(1+\alpha)}$ for $\alpha\leq 0.1$. If we conservatively assume an initial accretion velocity $v_{0}=0.01c_{s}$, where $c_{s}=\sqrt{{\partial p}/{\partial\rho}}$ is the sound speed derived from the equation of state (78), we obtain $v_{0}=0.01\alpha^{1/2}$.
@@
$F\left(t\right)=64\pi H_{A}^{2}M_{0}^{4}(1+\alpha)\int_{t_{0ac}}^{t}\rho_{ac}d\bar{t}=\frac{64\pi H_{A}M_{0}^{4}}{3}\left(\rho_{ac0}-\rho_{ac}\left(t\right)\right)$ (80)
@@
$M_{1}\left(t\right)=M_{0}+\frac{32\pi M^{3}_{0}}{3}\rho_{ac0}+18\pi\alpha M_{0}^{3}\rho_{ac}+\frac{64H_{A}M_{0}^{4}\pi}{3}\left(\rho_{ac0}-\rho_{ac}\left(t\right)\right),$ (85)
@@ 5 Conclusions
The formalism developed here is sufficiently general to be extended to a variety of related scenarios. In particular, a more detailed investigation of the backreaction effects on the fluid dynamics itself would be of considerable interest, especially concerning the structure of critical sonic points and the transonic nature of the flow in a cosmological background [28, 29, 44]. Moreover, the analysis may be generalized to other perfect fluids coupled to the expansion of the universe through alternative equations of state. Such extensions would allow one to assess the robustness of the qualitative features reported here.
@@
Acknowledgments
The work of L.F.R. is supported by FCT (Fundação para a Ciência e Tecnologia, Portugal) through the grant 2024.06413.BD. The work of O.B. is partially supported by FCT through the project 2024.00252.CERN. Both authors are partially supported by UID/04650 - Centro de Física das Universidades do Minho e do Porto. M. C. B. thanks the University of Porto for hosting part of this study.